### v1

@@ 3.3 Error Analysis
Trial Level: While the patients were enrolled in a specific trial, this did not preclude their eligibility for other concurrent trials. Consequently, the ranking numbers do not accurately reflect the model’s performance strength. To further assess this, we engaged a qualified clinical data expert to manually check the eligibility of patients for 10 randomly chosen trials that OncoLLM ranked in the top-1, but which were not the trials in which the patients were actually enrolled. We discovered that out of these 10 trials, the patients were deemed eligible for 9. Interestingly, the only trial for which a patient was ineligible was due to the presence of tumors in both breasts; the patient was eligible for the trial concerning the tumor in the left breast but not for the one in the right breast. This highlights OncoLLM’s effectiveness in identifying suitable trials for patients.
Criteria Level: To assess the capability of OncoLLM in generating accurate interpretation and citations of corresponding trial criterion to a patient, a set of questions (provided in supplement) and their corresponding responses generated by the model were randomly selected and reviewed by qualified medical professionals. These reviewers were tasked with verifying the correctness of the responses (categorized as ”Yes”, ”No”, or ”N/A”) and the accuracy of interpretation and citations provided with each answer. Our findings indicated that the accuracy of the final answers was 72.6%. When the final answer was deemed correct, the accompanying explanations were accurate 89.7% of the time. Additionally, citations included in the explanations were rated as correct 63.7% of the time and partially correct 8.5% of the time, highlighting OncoLLM’s effectiveness in delivering reliable information and its utility in supporting further manual verification processes.
@@ 4.3.2 Chunking and Retrieval Module
Each patient’s EHRs usually include hundreds of notes, which often surpass the context length capabilities of not only most open-source LLMs but also those of proprietary LLMs. This challenge requires the use of a semantic retrieval engine to manage and extract relevant information from each criterion [47]. We used the Python spaCy[48] tokenizer, configured with a one-sentence overlap, to chunk the patient notes efficiently. For the retrieval process, we utilized Open AI Ada embedding-based model and leveraged cosine similarity between the embeddings of the chunks and the queries derived from the inclusion and exclusion criteria of trials.
@@ 5 Discussion
Acknowledgements We are grateful for the assistance provided by medical professionals, doctors, and research coordinators who shared insights into existing clinical trial matching processes and challenges. We extend special thanks to Regina Schwind, Aniket Jaiswal, and Sebastien Rhodes for their help in interpreting results and offering valuable prompt engineering suggestions. Additionally, we express our gratitude to Sorena Nadaf, Warren Kibbe, and the entire Ci4CC (Cancer Informatics for Cancer Centers) community for facilitating this collaborative effort.
@@
S4  Prompts
This is the prompt used by these generator models- GPT3.5-Turbo, GPT4, Qwen14B-Chat and OncoLLM:
1[
2 {
3 "role" : "system",
4 "content" : {system prompt}
5 },
6 {
7 "role" : "user",
8 "content" : "
[…]
12
13 <CURRENT_DATE>
14 {current_date}
15 </CURRENT_DATE>
16
17 <PATIENT_INFO>
18 {patient_info}
19 </PATIENT_INFO>
20
21 <RELEVANT_CHUNKS>
22
23 CHUNK_ID_1: {retrieved_chunks[1][id]}
24 CHUNK_DATE_1: {retrieved_chunks[1][date]}
25 CHUNK_TITLE_1: {retrieved_chunks[1][title]}
26 CHUNK_CONTENT_1: {retrieved_chunks[1][text]}
27
28 CHUNK_ID_2: {retrieved_chunks[2][id]}
29 CHUNK_DATE_2: {retrieved_chunks[2][date]}
30 CHUNK_TITLE_2: {retrieved_chunks[2][title]}
31 CHUNK_CONTENT_2: {retrieved_chunks[2][text]}
32
33 ... (loop over all retrieved chunks)"
34 },
35 {
36 "role" : "assistant",
37 "content" : "‘‘‘json‘‘‘"
38 }
For Mistral-7B-Instruct and Mixtral-8x7B-Instruct models, since the option of a system message is unavailable, we split the system prompt (used for other models). The change is as following:
1[
2 {
3 "role" : "user",
4 "content" : {system prompt}
5 },
6 {
7 "role" : "assistant",
8 "content" : "I understood my role. Kindly provide the input data. I will make sure to provide the output in the provided JSON schema."
9 }
You are an Experienced Clinical Trial Matching Assistant, your task is to accurately answer clinical QUESTION about a cancer patient based on their Electronic Health Record (EHR) data. The QUESTION should be answered with a ’YES’, ’NO’, or ’N/A’ response. If the needed information is absent or irrelevant, your answer should be ’N/A’. To represent the degree of certainty in your answer, provide a numerical confidence score ranging from 1 (not confidence) to 5 (highly confident).
You will be provided with the PATIENT_INFO and RELEVANT_CHUNKS extracted from the patient’s EHR data to reason your answer for the QUESTION. Ensure that you are examining all the aspects of the patient documents and using medical reasoning to arrive at the final answer. Also you must ensure to consider the temporal aspects of the QUESTION as well, use the provided CURRENT_DATE as the current date for all your analysis. The success of your job depends on it, therefore take the necessary time to reason thoroughly.
In some cases, the RELEVANT_CHUNKS can have indirect information. For example, if there’s some information about the patient’s TNM staging according to AJCC (e.g., cT3N0M0 or T3, N0, M0), where T describes the size of the tumor and any spread of cancer into nearby tissue; N describes the spread of cancer to nearby lymph nodes; and M describes metastasis (spread of cancer to other parts of the body), then make sure to use these indirect information as well to form your answers. In some instances, if a patient’s condition is queried and the lab/test reports used to assess that condition are provided without indicating the presence of the disease in the patient, it is acceptable to presume that the patient does not have that condition.
In addition to your response, please cite the source of your information using the respective CHUNK_ID. Also, provide a detailed reasoning/explanation of your answer. Ensure your reasoning is logical, clear, succinct and medically correct.
## To ensure you answer the QUESTION very very accurately, you can follow the following step-by-step method to do the reasoning: 1. Understand the provided QUESTION properly, work out a strategy to answer a QUESTION like this. For some questions, temporal aspects is also important, take your time and make an informed decision whether to consider temporal aspects or not. 2. Understand the meaning and relationship between all the specified medical/clinical terms in the QUESTION and RELEVANT_CHUNKS. 3. Now work out a step-by-step logical deduction to answer the QUESTION verbally. For example: For the question ”Is the tumor size <10 cm?”, then work out like this: ”The patient has tumor size of 5 cm (CHUNK_ID) and 5 cm is less than 10 cm, which means tumor size <10 cm. For the question ”Does the patient have breast cancer?”, then work out like this: ”The patient’s primary site of cancer is Nipple (CHUNK_ID). Since Nipple is the primary site of breast cancer, which mean the patient has breast cancer. 4. Now based on your verbal answer decide the final answer from ’YES’, ’NO’, or ’N/A’. 5. Finally, provide a confidence score between 1-5, based on how confident you are.
Figure 14: System Prompt used for generating the Answer

### v2

@@ 3.3 Error Analysis
Trial Level: While the patients were enrolled in a specific trial, this did not preclude their eligibility for other concurrent trials. Consequently, the ranking numbers do not accurately reflect the model’s performance strength. To further assess this, we engaged a qualified clinical data expert to manually check the eligibility of patients for 10 randomly chosen trials that OncoLLM ranked in the top-1, but which were not the trials in which the patients were actually enrolled. We discovered that out of these 10 trials, the patients were deemed eligible for 9, effectively increasing the actual top-3 accuracy to $\approx 95\%$. Interestingly, the only trial for which a patient was ineligible was due to the presence of tumors in both breasts; the patient was eligible for the trial concerning the tumor in the left breast but not for the one in the right breast. This highlights OncoLLM’s effectiveness in identifying suitable trials for patients.
Criteria Level: To assess the capability of OncoLLM in generating accurate interpretation and citations of corresponding trial criterion to a patient, a set of questions (provided in supplement) and their corresponding responses generated by the model were randomly selected and reviewed by qualified medical professionals. These reviewers were tasked with verifying the correctness of the responses (categorized as ”Yes”, ”No”, or ”N/A”) and the accuracy of interpretation and citations provided with each answer. Our findings indicated that the accuracy of the final answers was 75.26%. When the final answer was deemed correct, the accompanying explanations were accurate 90.91% of the time. Additionally, citations included in the explanations were rated as correct 86.71% of the time and partially correct 6.29% of the time, highlighting OncoLLM’s effectiveness in delivering reliable information and its utility in supporting further manual verification processes.
@@ 4.3.2 Chunking and Retrieval Module
Each patient’s EHRs usually include hundreds of notes, which often surpass the context length capabilities of not only most open-source LLMs but also those of proprietary LLMs. This challenge requires the use of a semantic retrieval engine to manage and extract relevant information from each criterion [47]. We used the Python spaCy[48] tokenizer, configured with a one-sentence overlap, to chunk the patient notes efficiently. For the retrieval process, we utilized Azure OpenAI’s Ada embedding-based model and leveraged cosine similarity between the embeddings of the chunks and the queries derived from the inclusion and exclusion criteria of trials.
@@ 5 Discussion
Acknowledgements We are grateful for the assistance provided by medical professionals, doctors, and research coordinators who shared insights into existing clinical trial matching processes and challenges. We extend special thanks to Harsh Jain, Aniket Jaiswal, and Sebastien Rhodes for their help in interpreting results and offering valuable prompt engineering suggestions. Additionally, we express our gratitude to Sorena Nadaf, Warren Kibbe, and the entire Ci4CC (Cancer Informatics for Cancer Centers) community for facilitating this collaborative effort.